OpenStudy (mackenzie2013):

A gardener wants to create a rectangular vegetable garden in a backyard. She wants it to have a total area of 120 square feet, and it should be 12 feet longer than it is wide. What dimensions should she use for the vegetable garden? Round to the nearest hundredth of a foot. 4.95 feet by 16.95 feet 6.49 feet by 18.49 feet 10.95 feet by 22.95 feet 12.49 feet by 24.49 feet

OpenStudy (jiteshmeghwal9):

Let the width be w then length=w+12 Area=Length*width=w(w+12)=120 find w

OpenStudy (mackenzie2013):

I dont understand..

563blackghost (563blackghost):

We know that she wants the area to equal `120 feet`. We are told that the length is `12 feet more than the width` now turning that into an equation it would become `w+12`. And we don't know what `w` is....Now the way to find the area of a rectangle is using the equation `l x w = a` So we input what we know... \(\LARGE{l=w+12}\) \(\LARGE{a=120}\) \(\huge{(w+12)\times w =120}\)

OpenStudy (mackenzie2013):

@563blackghost , It would be B then.

563blackghost (563blackghost):

Correct :)